# Estimating the linear relation between variables that are never jointly observed

### v3

@@ 3.1 Moment estimators
If $\widehat{\boldsymbol{\mu}}_{1,X}$ is full rank, moment estimators of $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{d})$ can be built by considering the empirical counterpart of identification equations given in Lemma 2.1,
$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{Y}$ (3.1)
where $\widehat{\boldsymbol{\mu}}_{Y}=(\widehat{\mu}_{Y}^{1},\ldots,\widehat{\mu}_{Y}^{K})$. We consider in the following a slightly more general moment estimator of $\boldsymbol{\beta}$, by introducing a weight $w_{k}$ given to each group $k$ of observations, with $w_{k}>0$ and $\sum_{k=1}^{K}w_{k}=1$. The weighted moment estimator $\boldsymbol{\beta}$ is defined as the minimizer of
@@
$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}$ (3.2)
with $\mathbf{w}=(w_{1},\ldots,w_{K})$. The first moment estimator considered in (3.1) corresponds to equal weights $w_{k}=K^{-1}$.
@@
$\widehat{\sigma}_{\epsilon}^{2,M}$ $=\widehat{\sigma}_{Y}^{2}-(\widehat{\boldsymbol{\beta}}_{-0}^{M})^{\top}\widehat{\boldsymbol{\Gamma}}_{X}\widehat{\boldsymbol{\beta}}_{-0}^{M}.$ (3.3)
@@ 3.2 Optimal transport estimators
Estimators of $\boldsymbol{\beta}$ and $\sigma^{2}_{\epsilon}$ based on an optimal transport criterion are derived by minimizing the empirical version $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ of functional $\varphi(\boldsymbol{\gamma},\sigma^{2})$ defined by
$\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\widehat{\mu}_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}+\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (3.4)
Note that in absence of a priori information on the probability $\pi_{k}$ of observing group $k$, we can set $\pi_{k}=K^{-1}$. We denote by $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})$ minimizers of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ which are obtained with iterative optimization algorithms based on gradient descent. The algorithm can be initialized with $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}^{2,M})$
@@ 4 Consistency and asymptotic distribution
To study the asymptotic behavior of the estimators of $\boldsymbol{\beta}$ defined in previous section, we suppose that the number $K$ of groups is kept fixed and that for all groups and all variables $\mathbf{X}$ and $Y$, the number of observations tends to infinity. This means that $n_{\min}=\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})$ the smallest sample size among all experiments is also supposed to tend to infinity.
Under general moment conditions, without any particular assumption on the probability law of $Y$ and $\mathbf{X}$ we get in the following two Lemmas that the moment estimators and the Wasserstein distance based estimators are consistent. The proofs are postponed to the Appendix.
Lemma 4.1.
If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, and assumption $\mathbf{H}_{1}$ is fulfilled, the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}_{\epsilon}^{2,M})$ defined by (3.2) and (3.3) converges in probability to $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ when $n_{\min}$ tends to infinity.
@@
Lemma 4.2.
If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain 0, suppose that model (2.1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, then the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W}_{\epsilon})$ that minimize (3.4) converges in probability to $(\beta,\sigma^{2}_{\epsilon})$ when $n_{\min}$ tends to infinity.
As far as the asymptotic distribution of the estimators is concerned, and for sake of simplicity and lighter notations, we suppose from now on that the number of experiments is the same for all groups and all variables, that is to say $n=n_{y}^{1}=\cdots n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$.
Proposition 4.1.
Assume that the assumptions of Lemma 4.1 are fulfilled. Then as $n$ tends to infinity,
@@
This result is based on the CLT for empirical means and the application of the delta method, which involves computing the Jacobian of the inverse of matrices, making obtaining the explicit expression for $\boldsymbol{\Gamma}_{\beta_{M}}$ a difficult task when $d>1$.
Remark 4.1.
The weak convergence toward a Gaussian distribution presented in Proposition 4.1 would remain true, at the expense of heavier notations and a different asymptotic covariance matrix $\boldsymbol{\Gamma}_{\beta_{M}}$, provided that there exist two constants, $0<c\leq C$ such that
$0<c\leq\frac{\max(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}{\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}\leq C<+\infty,$ (4.1)
@@
The asymptotic normality of $\widehat{\boldsymbol{\beta}}^{W}$ relies on classical results for M-estimators recalled in the Appendix (see Theorem A.4). Note that since we estimate simultaneously $\beta$ and $\sigma_{\epsilon}^{2}$, an additional condition on the existence of the moments of order four for the covariates is required.
Proposition 4.2.
If model (2.1) holds and hypothesis $\mathbf{H}_{1}$ is fulfilled, ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{4})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain $(0,0)$, then, as $n$ tends to infinity,
@@ 5 The particular case of simple linear regression
To show the difficulty, let us examine the case of simple linear regression, that is to say $d=1$. The following linear model
@@
is supposed to hold and if there are two groups $k$ and $j$ such that $\mu_{X}^{k}\neq\mu_{X}^{j}$, the identification assumption $\mathbf{H}_{1}$ is fulfilled and $\beta_{0}$ and $\beta_{1}$ can be uniquely determined. The moment estimators of $\beta_{0}$ and $\beta_{1}$ defined in (3.2) have simple expressions,
$\widehat{\beta}_{0}$ $=\widehat{\mu}_{Y,w}-\widehat{\beta}_{1}\widehat{\mu}_{X,w}$ (5.1)
$\widehat{\beta}_{1}$ $=\frac{\sum_{k=1}^{K}w_{k}\widehat{\mu}_{X}^{k}\widehat{\mu}_{Y}^{k}-\widehat{\mu}_{X,w}\widehat{\mu}_{Y,w}}{\sum_{k=1}^{K}w_{k}\left(\widehat{\mu}_{X}^{k}\right)^{2}-\widehat{\mu}_{X,w}^{2}}$ (5.2)
@@
We focus on the asymptotic variance of the estimator $\widehat{\beta}_{1}$ of the slope parameter $\beta_{1}$, which is often the parameter of interest. We have
Lemma 5.1.
Suppose that the model (2.1) is true, with $d=1$ and $K\geq 2$. If there are two groups $k$ and $j$ such that $\mu_{X}^{k}\neq\mu_{X}^{j}$ the vector $(\beta_{0},\beta_{1})$ is identifiable and, as $n$ tends to infinity,
@@
Lemma 5.1 shows us that even in a very simple framework (only one covariate), the asymptotic variance of the estimator of the slope $\beta_{1}$ is quite complicated. It also reveals that minimizing the asymptotic variance $\widehat{\beta}_{1}$ with respect to the weights $w_{k}$ is not a simple task.
@@ 6 Bootstrapping for confidence intervals
Since, as noted in the previous section, it is complicated to compute explicitly the asymptotic variance matrix of $\widehat{\boldsymbol{\beta}}^{M}$ and $\widehat{\boldsymbol{\beta}}^{W}$, we consider stratified bootstrap approaches in order to build confidence sets for $\boldsymbol{\beta}$.
@@
Our bootstrap procedure takes account of the independence between the different groups $k=1,\ldots,K$, as well as the independence of inputs $(X_{1}^{k},\ldots,X_{d}^{k})$ and output $Y^{k}$ within each group, meaning more formally that, given $G=k$, the joint probability measure $\mathbb{P}^{k}$ of $Y$ and $\mathbf{X}$ is a product measure of the marginal measures $\mathbb{P}^{k}=\mathbb{P}_{Y}^{k}\otimes\mathbb{P}_{\mathbf{X}}^{k}$.
@@
Within each group $k$, we draw, with equal probability and with replacement, $n_{y}^{k}$ observations among $Y_{1}^{k},\ldots,Y_{n_{y}^{k}}^{k}$, We also draw independently, with equal probability and with replacement, $n_{x}^{k}$ observations among $\mathbf{X}_{1}^{k},\ldots,\mathbf{X}_{n_{x}^{k}}^{k}$. We denote by $\mu_{Y}^{k*}$ and by $\boldsymbol{\mu}_{X}^{k*}$ the empirical means and by $\sigma_{Y,k}^{2,*}$ and by $\boldsymbol{\Gamma}_{X}^{k*}$ the empirical variances of $Y$ and $\mathbf{X}$ in these bootstrap samples. Bootstrapped estimators $\boldsymbol{\beta}^{M,*}$ and $\boldsymbol{\beta}^{W,*}$ of $\boldsymbol{\beta}$ can now be computed by replacing the empirical moments by the bootstrap moments in (3.2) and (3.4).
To build confidence sets for the components of $\boldsymbol{\beta}$ based on previous bootstrap procedure, the bootstrap percentile technique, described in Chapter 4 of Shao and Tu, (1995), is simple to use.
@@
It can be noted that our estimators are smooth functions of sample means so that classical bootstrap theory applies (see for example Shao and Tu, (1995), Chapter 3). For simplicity, as in Proposition 4.1, we suppose that $n=n_{y}^{1}=\ldots=n_{y}^{K}=n_{x}^{1}=\ldots=n_{x}^{K}$. Because of the considered experimental design, our global "empirical distribution" is made of products of marginal empirical distributions, the bootstrap for means is almost surely consistent for the Kolmogorov metric, and with Theorem 3.1 in Shao and Tu, (1995) the same result holds for the estimators of $\boldsymbol{\beta}$ considered in this work. Then, the application of Theorem 4.1 in Shao and Tu, (1995) allows to conclude that bootstrap percentile method gives consistent confidence sets for each component of $\boldsymbol{\beta}$.
Proposition 6.1.
@@
Suppose that $\mathbb{E}(Y^{2})<\infty$ and $\mathbb{E}\|\mathbf{X}\|^{2}<\infty$ and hypothesis $\mathbf{H}_{1}$ is fulfilled. Then as $n\to+\infty$, the bootstrap estimator $\boldsymbol{\beta}^{M,*}$ is strongly consistent for $\boldsymbol{\beta}$ in the Kolmogorov metric. Consider a risk $\alpha\in(0,1)$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides, for a given nominal level $1-\alpha$, a consistent confidence set.
@@
We can also state a similar result for the estimators minimizing the Wasserstein distance, under slightly more restrictive moment conditions and compactness assumption.
Proposition 6.2.
@@
Suppose that the assumptions of Proposition 4.2 are fulfilled. As $n\to+\infty$, the bootstrap estimator $(\boldsymbol{\beta}^{W,*},\sigma_{\epsilon}^{2,W,*})$ is strongly consistent for $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ in the Kolmogorov metric. Consider a risk $\alpha\in(0,1)$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides, for a given nominal level $1-\alpha$, a consistent confidence set.
@@
Propositions 6.1 and 6.2 are very important for practical applications since they ensure that even if we are not able to compute the explicit expression of the asymptotic variance of our estimators, we can still construct asymptotic confidence intervals thanks to simple bootstrap procedures.

### v4

@@ 3.1 Moment estimators
If $\widehat{\boldsymbol{\mu}}_{1,X}$ is full rank, moment estimators of $\boldsymbol{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{d})$ can be built by considering the empirical counterpart of the identification equations given in Lemma 2.1,
$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\widehat{\boldsymbol{\mu}}_{Y}$ (4)
where $\widehat{\boldsymbol{\mu}}_{Y}=(\widehat{\mu}_{Y}^{1},\ldots,\widehat{\mu}_{Y}^{K})$. In the following, we consider a slightly more general moment estimator of $\boldsymbol{\beta}$, by introducing a weight $w_{k}$ given to each group $k$ of observations, with $w_{k}>0$ and $\sum_{k=1}^{K}w_{k}=1$. The weighted moment estimator $\boldsymbol{\beta}$ is defined as the minimizer of
@@
$\widehat{\boldsymbol{\beta}}^{M}$ $=\left(\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{1,X}\right)^{-1}\widehat{\boldsymbol{\mu}}_{1,X}^{\top}\mathbf{w}\widehat{\boldsymbol{\mu}}_{Y}$ (5)
where $\mathbf{w}$ is a diagonal matrix with diagonal elements $w_{1},\ldots,w_{K}$. The first moment estimator considered in (4) corresponds to the case with equal weights $w_{k}=K^{-1}$.
@@
$\widehat{\sigma}_{\epsilon}^{2,M}$ $=\widehat{\sigma}_{Y}^{2}-(\widehat{\boldsymbol{\beta}}_{-0}^{M})^{\top}\widehat{\boldsymbol{\Gamma}}_{X}\widehat{\boldsymbol{\beta}}_{-0}^{M}.$ (6)
@@ 3.2 Optimal transport estimators
Estimators of $\boldsymbol{\beta}$ and $\sigma^{2}_{\epsilon}$ based on an optimal transport criterion are derived by minimizing the empirical version $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ of a functional $\varphi(\boldsymbol{\gamma},\sigma^{2})$ defined by
$\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ $=\sum_{k=1}^{K}\pi_{k}\left[(\widehat{\mu}_{Y}^{k}-\gamma_{0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k})^{2}+\left(\widehat{\sigma}_{Y,k}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma^{2}}\right)^{2}\right].$ (7)
Note that in absence of a priori information on the probability $\pi_{k}$ of observing group $k$, we can set $\pi_{k}=K^{-1}$. We denote by $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})$ the minimizers of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$, which are obtained with iterative optimization algorithms based on gradient descent. The algorithm can be initialized with $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}^{2,M})$
@@ 4 Consistency and asymptotic distribution
To study the asymptotic behavior of the estimators of $\boldsymbol{\beta}$ defined in the previous section, we assume that the number $K$ of groups is kept fixed, and that for all groups and all variables $\mathbf{X}$ and $Y$, the number of observations tends to infinity. This means that $n_{\min}=\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})$, the smallest sample size among all experiments, should also tend to infinity.
Lemma 4.1
If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, and the assumption $\mathbf{H}_{1}$ is fulfilled, the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{M},\widehat{\sigma}_{\epsilon}^{2,M})$ defined by (5) and (6) converges in probability to $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ when $n_{\min}$ tends to infinity.
@@
Lemma 4.2
If ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{2})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain 0, if the model (1) holds and the hypothesis $\mathbf{H}_{1}$ is fulfilled, then the sequence of estimators $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W}_{\epsilon})$ that minimize (7) converges in probability to $(\beta,\sigma^{2}_{\epsilon})$ when $n_{\min}$ tends to infinity.
As far as the asymptotic distribution of the estimators is concerned, and for the sake of simplicity and simpler notation, from now on we will assume that the number of experiments is the same for all groups and all variables, i.e. $n=n_{y}^{1}=\ldots=n_{y}^{K}=n_{x}^{1}=\ldots=n_{x}^{K}$.
Proposition 4.1
If the assumptions of Lemma 4.1 are fulfilled, as $n$ tends to infinity,
@@
This result is based on the central limit theorem for empirical means and the application of the delta method, which involves computing the Jacobian of the inverse of matrices, making it difficult to obtain the explicit expression for $\boldsymbol{\Gamma}_{\beta_{M}}$ when $d>1$.
Remark 4.1
The weak convergence toward a Gaussian distribution presented in Proposition 4.1 remains true, at the expense of heavier notation and a different asymptotic covariance matrix $\boldsymbol{\Gamma}_{\beta_{M}}$, provided that there exist two constants, $0<c\leq C$ such that
$0<c\leq\frac{\max(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}{\min(n_{y}^{1},\cdots,n_{y}^{K},n_{x}^{1},\cdots,n_{x}^{K})}\leq C<+\infty,$ (8)
@@
The asymptotic normality of $\widehat{\boldsymbol{\beta}}^{W}$ relies on classical results for M-estimators recalled in Supplementary Material (see Theorem A.4). Note that since we estimate $\beta$ and $\sigma_{\epsilon}^{2}$ simultaneously, an additional condition on the existence of the moments of order four is required for the covariates.
Proposition 4.2
If the model (1) holds, the hypothesis $\mathbf{H}_{1}$ is fulfilled, ${\mathbb{E}}(Y^{2})<+\infty$ and ${\mathbb{E}}(\|\mathbf{X}\|^{4})<+\infty$, $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ and $\Theta$ is a compact set that does not contain $(0,0)$, then, as $n$ tends to infinity,
@@ 5 The particular case of simple linear regression
To illustrate the difficulty, consider the case of a simple linear regression, i.e. $d=1$. The following linear model
@@
is assumed to hold, and if there are two groups $k$ and $j$ such that $\mu_{X}^{k}\neq\mu_{X}^{j}$, the identification assumption $\mathbf{H}_{1}$ is fulfilled and $\beta_{0}$ and $\beta_{1}$ can be uniquely determined. The moment estimators of $\beta_{0}$ and $\beta_{1}$ defined in (5) have simple expressions:
$\widehat{\beta}_{0}$ $=\widehat{\mu}_{Y,w}-\widehat{\beta}_{1}\widehat{\mu}_{X,w}$ (9)
$\widehat{\beta}_{1}$ $=\frac{\sum_{k=1}^{K}w_{k}\widehat{\mu}_{X}^{k}\widehat{\mu}_{Y}^{k}-\widehat{\mu}_{X,w}\widehat{\mu}_{Y,w}}{\sum_{k=1}^{K}w_{k}\left(\widehat{\mu}_{X}^{k}\right)^{2}-\widehat{\mu}_{X,w}^{2}}$ (10)
@@
We focus on the asymptotic variance of the estimator $\widehat{\beta}_{1}$ of the slope parameter $\beta_{1}$, which is often the parameter of interest.
Lemma 5.1
Suppose that the model (1) is true, with $d=1$ and $K\geq 2$. If there are two groups $k$ and $j$ such that $\mu_{X}^{k}\neq\mu_{X}^{j}$ the vector $(\beta_{0},\beta_{1})$ is identifiable and, as $n$ tends to infinity,
@@
Lemma 5.1 shows that even in a very simple framework (only one covariate) the asymptotic variance of the estimator of the slope $\beta_{1}$ is quite complicated. It also reveals that minimizing the asymptotic variance $\widehat{\beta}_{1}$ with respect to the weights $w_{k}$ is not a simple task.
@@ 6 Bootstrapping for confidence intervals
Since, as noted in the previous section, it is complicated to explicitly compute the asymptotic variance matrix of $\widehat{\boldsymbol{\beta}}^{M}$ and $\widehat{\boldsymbol{\beta}}^{W}$, we consider stratified bootstrap approaches in order to build confidence sets for $\boldsymbol{\beta}$. Our bootstrap procedure takes into account the independence between the different groups $k=1,\ldots,K$ as well as the independence of the inputs $(X_{1}^{k},\ldots,X_{d}^{k})$ and the output $Y^{k}$ within each group. More formally, given $G=k$, the joint probability measure $\mathbb{P}^{k}$ of $Y$ and $\mathbf{X}$ is a product measure of the marginal measures $\mathbb{P}^{k}=\mathbb{P}_{Y}^{k}\otimes\mathbb{P}_{\mathbf{X}}^{k}$.
@@
Within each group $k$, we draw, with equal probability and with replacement, $n_{y}^{k}$ observations among $Y_{1}^{k},\ldots,Y_{n_{y}^{k}}^{k}$. We also draw independently, with equal probability and with replacement, $n_{x}^{k}$ observations among $\mathbf{X}_{1}^{k},\ldots,\mathbf{X}_{n_{x}^{k}}^{k}$. We denote by $\mu_{Y}^{k*}$ and by $\boldsymbol{\mu}_{X}^{k*}$ the empirical means and by $\sigma_{Y,k}^{2,*}$ and by $\boldsymbol{\Gamma}_{X}^{k*}$ the empirical variances of $Y$ and $\mathbf{X}$ in these bootstrap samples. Bootstrapped estimators $\boldsymbol{\beta}^{M,*}$ and $\boldsymbol{\beta}^{W,*}$ of $\boldsymbol{\beta}$ can now be computed by replacing the empirical moments by the bootstrap moments in (5) and (7). To build confidence sets for the components of $\boldsymbol{\beta}$ based on this bootstrap procedure, the bootstrap percentile technique described in Chapter 4 of Shao and Tu, (1995) can be applied.
@@
It can be noted that our estimators are smooth functions of the sample means, so classical bootstrap theory applies (see for example Shao and Tu, (1995), Chapter 3). For simplicity, we assume, as in Proposition 4.1, that $n=n_{y}^{1}=\ldots=n_{y}^{K}=n_{x}^{1}=\ldots=n_{x}^{K}$. Because of the experimental design considered, our global ”empirical distribution” consists of products of marginal empirical distributions, the bootstrap for the means is almost surely consistent for the Kolmogorov metric, and with Theorem 3.1 in Shao and Tu, (1995) the same result holds for the estimators of $\boldsymbol{\beta}$ considered in this work. The application of Theorem 4.1 in Shao and Tu, (1995) allows to conclude that the bootstrap percentile method gives consistent confidence bounds for each component of $\boldsymbol{\beta}$.
Proposition 6.1
@@
Suppose that $\mathbb{E}(Y^{2})<\infty$ and $\mathbb{E}\|\mathbf{X}\|^{2}<\infty$ and the hypothesis $\mathbf{H}_{1}$ is fulfilled. Then as $n\to+\infty$, the bootstrap estimator $\boldsymbol{\beta}^{M,*}$ is strongly consistent for $\boldsymbol{\beta}$ in the Kolmogorov metric. Consider a risk $\alpha\in(0,1)$, then for a given nominal level $1-\alpha$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides a consistent confidence set.
@@
We can also state a similar result for the estimators minimizing the Wasserstein distance, under slightly more restrictive moment conditions and a compactness assumption.
Proposition 6.2
@@
Suppose that the assumptions of Proposition 4.2 are fulfilled. As $n\to+\infty$, the bootstrap estimator $(\boldsymbol{\beta}^{W,*},\sigma_{\epsilon}^{2,W,*})$ is strongly consistent for $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ in the Kolmogorov metric. Consider a risk $\alpha\in(0,1)$, then for a given nominal level $1-\alpha$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides a consistent confidence set.
@@
Propositions 6.1 and 6.2 are very important for practical applications since they ensure that even if we are unable to compute the explicit expression of the asymptotic variance of our estimators, the asymptotic confidence intervals can still be constructed with a simple bootstrap procedure.